OpenStudy (anonymous):

5/-8 + -6/5???? HELP!!

OpenStudy (dumbcow):

find the common denominator...just like you've done before

OpenStudy (anonymous):

i dont know how to find the common denominator

OpenStudy (anonymous):

:(

OpenStudy (dumbcow):

its the LCM of 8 and 5

OpenStudy (anonymous):

how do u find the LCM?

OpenStudy (dumbcow):

8+8+8+8+8 5+5+5+5+5 when do they meet?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

??

OpenStudy (anonymous):

now what do i do with it??

OpenStudy (dumbcow):

8+8=16+8=24+8=32+8=40 5+5=10+5=15+5=20+5=25+5=30+5=40 they meet at 40 or 8*5 that is now our new denominator, but we have to change numerators so they are equivalent -5/8 = -25/40 i multiplied both top and bottom by 5 -6/5 = -48/40 multiply top and bottom by 8

OpenStudy (dumbcow):

once you have the new fractions with same denominator just add the top numbers -25 + -48 = -73 = -73/40

OpenStudy (anonymous):

ohhh okk

OpenStudy (anonymous):

that was right! ok last one!

OpenStudy (dumbcow):

if you have to just multiply the bottom numbers, that will always give you a common denominator

OpenStudy (dumbcow):

always look for 2 neg signs next to each other, they cancel each out do this first

OpenStudy (dumbcow):

do you want an explanation or just the answer

OpenStudy (dumbcow):

61/18